# Large Language Models as Proxies for Theories of Human Linguistic Cognition

### v1

@@ 4.3 Test case: full-reverse
We find that the perturbed versions of Italian, Russian and Hebrew are projected as easier than attested English (see Fig. 4), even though the perturbation is considered humanly impossible. In this case, ease of learning according to Kallini et al. (2024)’s perplexity metric does not provide support for the idea that the target theory can explain away the typological asymmetry through asymmetries of ease of learning. As mentioned above, this conclusion is weak. It is possible that the target theory is the best explanation and that the typology is explained with the help of factors other than ease of learning. It is also very possible that Kallini et al.’s perplexity metric is a poor way to evaluate ease of approximation by the LLM. All we can conclude at present is that proponents of the Proxy View have work to do in light of the LLM’s performance in this case, and that previous results only provide a limited picture.
Figure 4: Validation perplexity during training for attested (baseline) and full-reverse versions of English, Russian, Italian, and Hebrew.
@@ 4.4 Test case: switch-indices
We turn to another asymmetry that the Proxy View does not at present explain. Consider $\Pi=$ switch-indices, where the tokens at index 0 and index 2 in every sentence are switched:
@@
If the target theory can account for the typological asymmetry, this is again not detected through the perplexity curve of the LLM proxy: each language is wrongly projected as harder than its perturbed version, and the perplexity delta between each two attested languages is significantly larger (see Fig. 5). In all cases, there is an unattested language that is projected as easier than an attested language.
Figure 5: Validation perplexity during training for the attested (baseline) and switch-indices versions of English, Italian, Hebrew and Russian.
@@ 4.5 Test case: token-hop, no-hop
Since additional textual material is inserted, in a way that might confound perplexities over time, we compare the perplexities achieved by our model on token-hop datasets to no-hop datasets, in which the marker token is inserted exactly after each verb. This achieves control for the effect of the marker token itself, such that difference in perplexities can be attributed solely to the effect of the token-counting generalization. The no-hop version adheres to attested generalizations, as it can be thought of as post-verbal clitic that marks for parts of speech. Learning token-hop on the other hand crucially requires counting tokens, an ability that is particularly non-humanlike, both because human languages have no atomic concept that corresponds to tokens, and because counting rules violate the strictly hierarchal nature of linguistic generalizations across languages. Although the distance of the token from its associated verb has an effect on perplexity, we find that the no-hop version of English is still harder to learn than token-hop Italian and Russian (see Fig. 6). Unperturbed English datasets are consistently projected as harder than token-hop perturbed versions of Italian and Russian, despite a clear typological skew against counting rules. The same point is made with the validation perplexities of token-hop Italian and no-hop Russian patterning very similarly. The best explanation of linguistic typology should predict a clear divide between unattested languages that crucially require counting linguistic elements and attested languages. Ease-of-learning via LLM does not explain this clear typological skew.
Figure 6: Validation perplexity during training for the no-hop (baseline) and token-hop versions of English, Italian, and Russian.
@@ 5 Conclusion
LLMs have been presented as a challenge to more traditional approaches to the study of HLC, and in particular to generative linguistics. A blunt version of this idea is the LLM Theory, promoted by Piantadosi (2023), which maintains that current LLMs are themselves good theories of HLC. Current LLMs can, of course, be viewed as theories of HLC, but this treatment does these models no favors, as the literature was quick to note.
A more cautious approach is the Proxy View, promoted by Wilcox et al. (2023) and Mahowald et al. (2024), which maintains that current LLMs are good proxies - as far as various behavioral measures are concerned - for good theories of HLC that are more linguistically-neutral than generative theories. No commitment to shared essentials between LLMs and the relevant good theories is implied. The Proxy View is of course perfectly coherent, but differently from the LLM Theory it is too vague about the theories of HLC that it champions to be of much direct use within empirical science. Empirical science is based on inference to the best explanation and proceeds through competition between reasonably well-understood theories, and saying just that one’s favorite theory learns roughly as well as an LLM is insufficient for this purpose. One would like to see enough detail at least for a proper comparison of the relevant theory to generative linguistics in light of the discoveries and considerations in the literature of the past 70 years or so. More than anything else, the present paper is a plea for supporters of the Proxy View to provide this kind of detail and evaluation for the theory they have in mind.
Still, already at this stage we can outline an empirical evaluation of the Proxy View on various kinds of test cases, and here we focused on two: alignment with the stimulus and cross-linguistic variation. The results were not encouraging for the Proxy View. The LLMs failed to even approximate key pieces of knowledge that humans have, even when the models were trained on much larger corpora than children are exposed to. The LLMs also had an easier time approximating various typologically-unattested languages than actual human languages, at least with respect to a success criterion used in Kallini et al. (2024).
Appendix A Section 3 - Appendix
@@ A.1 General Notes
In all experiments, probabilities are evaluated on the critical region in boldface (as in the example sentences in Tables A.2, A.2, A.2). For models listed as unidirectional, the probability value takes into consideration the left context of the evaluated token, whereas for bidirectional models, the probability value factors in both the left and the right context. The bidirectional models we consider were pretrained with masked language modeling (fill-mask).
In all generated sentences, the critical region unambiguously disambiguates the sentence as an instance of the phenomenon under consideration (there is no other grammatical reading). To promote lexical variation in target regions, possible target tokens in the CFGs were chosen at random from a POS-tagged English Wikipedia dump. We filtered out potential target tokens that were not part of the vocabulary of one the models considered in our setup, which are detailed in Table 5.
@@
CHILDES Transformer 20M Uni Yedetore et al. 2023
BabyLM 10M (strict-small) 10M Bi Warstadt et al. 2023b
Wikipedia Transformer 15M Uni Trained by authors
BabyLM 100M 100M Bi Warstadt et al. 2023b
BERT Base Uncased $\approx$3.5B Bi Devlin et al. 2018b
GPT2 $\approx$8B Uni Radford et al. 2019b
Llama 3.2-3b $\approx$9T Uni Touvron et al. 2023
Table 5: Models used in the experiments in Section 3.2.
@@ A.2 ATB and PG
For the ATB and PG experiments, we replicate the experimental setup from Lan et al. 2024b with a wider variety of test sentences to increase robustness. We also unify the templates for both phenomena so that they use the same lexical choices where relevant (e.g., proper names). From each template, represented by a context free grammar, we generated tuples of sentences that correspond to the (+filler,+gap), (+filler,-gap) conditions (see CFGs and example sentences in Tables A.2 and A.2). We then evaluated the success criterion $P(+filler,-gap)<P(+filler,+gap)$ on a sample of 10,000 tuples for each phenomenon. For ATB, sentence structure was crucially changed such that the two conjuncts do not agree in number, to prevent a problematic non-ATB reading present in some of the Lan et al. test sentences. Underlined words alternate according to the ±$\pm filler$ condition; words in bold mark the position where the ±$\pm gap$ condition becomes evident and probability is measured. We evaluate the performance of four models from Table 5. We do not consider bidirectional models for the PG and ATB experiments because they are highly sensitive to whether the token is the last token in the sentence, which in the ATB and PG case is one of the minimal differences between the conditions we consider.
@@
$S\rightarrow\langle{PREAMBLE}\rangle\ \langle{\pm F}\rangle\ \langle{LINK}\rangle\ \langle{\pm G}\rangle$
@@
$\langle{+F}\rangle\ \rightarrow\ \underline{\textit{{which}}\ \langle{PLNOUN}\rangle\ }\langle{VP1}\rangle\ \ \langle{ADV1}\rangle$
[…]
$\langle{-G}\rangle\ \rightarrow\langle{LINK}\rangle\ \ \langle{VP2}\rangle\ \ \langle{\textbf{OBJ}}\rangle\ \ \langle{ADV2}\rangle$
@@
$\langle{VP1}\rangle\ \rightarrow\langle{VP1\_SIMPLE}\rangle\ \ |\ \langle{VP1\_COMPLEX}\rangle$
[…]
$\langle{VP2}\rangle\ \rightarrow\langle{VP2\_SIMPLE}\rangle\ \ |\ \langle{VP2\_COMPLEX}\rangle$
[…]
@@
$\Rightarrow$ I know which boys John met recently and is going to hug soon. (+filler,+gap)
$\Rightarrow$ *I know which boys John met recently and is going to hug you soon. (+filler,-gap)
Table 6: Excerpt from the context-free grammar used to generate Across-the-Board sentences for the experiments in Section 3.
Table 7: Excerpt from the context-free grammar used to generate Parasitic Gap sentences for the experiments in Section 3.
@@ A.3 TTE
To create test sets, we generate 1,178,496 quadruplets, out of which we randomly sample 10,000 quadruplets per experiment. Paradigms are varied lexically in several ways: we first create paradigms that differ structurally from each other, to control for the possibility that our results are an artifact of the specific sentence structure introduced in the example sentences in Table A.2. We also introduce lexical variation to all non-target elements, such as the main subject, the embedding verb, the wh-element and the verb inside the wh-clause. Most importantly, lexical choices for target verbs and nouns are varied using 30 target nouns and 30 target verbs. We perform five inference experiments on five such samples and consider average accuracy over all experiments. Quadruplets are evaluated with the two success criteria introduced in Section 3.3.
Table 8: Excerpt from the context-free grammar used to generate That-Trace-Effect sentences for the experiments in Section 3.
Appendix B Section 4 - Appendix

### v2

@@ 3.2 Test case: partial-reverse
Reversing linguistic data is generally considered an effective way to generate unnatural data, since it disrupts the language’s consistent word orders (Mitchell and Bowers, 2020), and in the case of partial-reverse, also unnaturally entangles dependencies involving constituents that cross the <rev> token boundary. The target, relatively linguistically-neutral theory $H_{3}$ in this case might be a version of a generative theory $H_{1}$ in which there is a parameter for reverse. Any grammar that is statable within $H_{3}$ now comes in two variants. If the parameter is set to 0, the grammar works as in the original generative theory. But if the parameter is set to 1, every derivation ends with reversing the output starting from a random point. This may or may not be the target theory that Kallini et al. have in mind; other target theories are of course imaginable, and Kallini et al. do not comment explicitly on this. We note that, like Kallini et al.’s other perturbations, reversal seems quite far away from the typological questions that generative linguistics have usually concerned themselves with. Here we stay with Kallini et al.’s perturbations and setup simply in order to illustrate the relevant methodological considerations, but a proper exploration of the Proxy View would need to look at those cross-linguistic patterns that have informed linguistic research.
Рис. 3: Validation perplexity during training for English, Italian, and Russian and their partial-reverse perturbations. The results indicate that the learning curves for attested languages are consistently below those of their partial-reverse counterparts.
We find that validation perplexities follow an attested-unattested divide within languages: the partial-reverse perturbed versions of English, Italian and Russian are harder to learn than their attested counterparts (see Figure 3). [^14] The partial-reverse languages are compared to a modified baseline of English in which the marker <rev> is inserted randomly without reversing, to control for the effect for additional textual material in the sentence on perplexities. Since we use the same seed for both perturbations, the markers are inserted in the same indices (see Example 3.2).
This, as discussed above, could be part of an argument that linguistic biases are not needed to rule out partial-reverse languages. Consider an $H_{3}$ that is a serious contender as a theory of HLC, addressing the various arguments regarding HLC in the literature. Suppose that $H_{3}$ is relatively linguistically neutral in the sense of allowing not just the languages sanctioned by a standard generative $H_{1}$ but also their partial-reverse versions. An argument in favour of $H_{1}$ and against $H_{3}$ might be that these partial-reverse variants are in fact unattested. Our results, following Kallini et al. 2024, can be taken as a response to this objection: using GPT-2 as a proxy for $H_{3}$, there is preliminary reason to think that attested languages might be slightly easier to learn than their partial-reverse variants, which might be part of an explanation of why the latter are unattested. This, in turn, could allow us to keep $H_{3}$ as a contender.
Other perturbations paint a different picture, as we now show.
3.3 Test case: full-reverse
We now consider the perturbation full-reverse from Kallini et al. 2024, where each sentence is reversed in its entirety and a special marker token is randomly inserted:
[…]
We find that across languages, the perturbed versions of Italian, Russian and Hebrew are projected as easier than attested English (see Figure 4), even though the perturbation is not considered humanly impossible. Within languages, Italian is projected as harder than its perturbed version. In this case, ease of learning according to Kallini et al. 2024’s perplexity metric does not provide support for the idea that the target theory can explain away the typological pattern through asymmetries of ease of learning.
Рис. 4: Validation perplexity during training for attested (baseline) and full-reverse versions of English, Russian, Italian, and Hebrew.
@@ 3.4 Test case: switch-indices
We turn to another typological asymmetry that the Proxy View does not at present explain. Consider switch-indices, where the tokens at index 0 and index 2 in every sentence are switched:
@@
If the target theory can account for the typological asymmetry, this is again not detected through the perplexity curve of the LLM proxy: within languages, each attested language is wrongly projected as harder than its perturbed version (see Figure 5).
Рис. 5: Validation perplexity during training for the attested (baseline) and switch-indices versions of English, Italian, Hebrew and Russian.
@@ 3.5 Test case: token-hop, no-hop
We look at one final typological asymmetry, based on the token-hop perturbation employed by Kallini et al. (2024). For each of English, Italian, and Russian, we create a perturbed variant by inserting a new marker token three tokens after each verb: [^15] New marker tokens were randomly generated as single-character tokens that do not already exist in the model’s vocabulary (English: “v”, Italian: “v”, Russian: “Ч”).
@@
[…]
Since additional textual material is inserted, in a way that might confound perplexities over time, we compare the perplexities achieved by our model on token-hop datasets to no-hop datasets, in which the marker token is inserted exactly after each verb. This controls for the effect of the marker token itself, such that difference in perplexities can be attributed solely to the effect of the token-counting generalization. The no-hop version adheres to attested generalizations, as it can be thought of as a post-verbal clitic that marks for parts of speech. Learning token-hop on the other hand crucially requires counting tokens, an ability that is particularly non-humanlike, because rules that require counting linear positions in a string violate the strictly hierarchical nature of linguistic generalizations across languages. Although within languages the distance of the token from its associated verb has an effect on perplexity, across languages we find that the no-hop version of English is still harder to learn than token-hop Italian and Russian (see Figure 6). Unperturbed English datasets are consistently projected as harder than token-hop perturbed versions of Italian and Russian, despite a clear typological skew against counting rules. The same point is made with the validation perplexities of token-hop Italian and no-hop Russian patterning very similarly.
Рис. 6: Validation perplexity during training for the no-hop (baseline) and token-hop versions of English, Italian, and Russian.
3.6 Limitations of the typology experiment
In this section, we have considered ease of learning by an LLM as a proxy for a linguistically-neutral theory that can explain typological patterns. It should be emphasized that the inferences from this argument are particularly weak. On the one hand, even if the linguistically-neutral theory $H_{3}$ derives a learning asymmetry between pattern-satisfying and pattern-defying languages, this asymmetry might not suffice to account for a robust typological pattern. On the other hand, even if the linguistically-neutral theory $H_{3}$ does not derive a learning asymmetry (as was the case with respect to several perturbations above), it might still be able to account for it in terms of other factors, such as communicative pressure. In order to overcome this, one could try to situate the learning component within a model of cultural evolution, which allows for the amplification of small asymmetries over generations (see Kirby et al., 2007, Niyogi and Berwick, 2009, and Brochhagen et al., 2018, among others) and can potentially balance multiple interacting pressures. We believe that such a model is the proper context for the use of ease of learning by LLMs to evaluate typological asymmetries, but following Kallini et al., whose setup we use here, we leave such a model for future work and only assess currently available models. All we can conclude at present is that, with the possible exception of partial reverse, Kallini et al.’s setup does not currently provide support to the undisclosed $H_{3}$ on the Proxy View. [^16] For further discussion of Kallini et al. 2024 and its limitations see Hunter 2025.
@@ 4 Conclusion
LLMs have been presented as a challenge to more traditional approaches to the study of HLC, and in particular to generative linguistics. A blunt version of this idea is the LLM Theory, promoted by Piantadosi (2023), which maintains that current LLMs are themselves good theories of HLC. Current LLMs can, of course, be viewed as theories of HLC, but this treatment does these models no favours, as the literature was quick to note.
A more cautious approach is the Proxy View, possibly the position of Wilcox et al. (2023) and Mahowald et al. (2024), which maintains that current LLMs are good proxies – as far as various behavioural measures are concerned – for good theories of HLC that are more linguistically-neutral than generative theories. No commitment to shared essentials between LLMs and the relevant good theories is implied. The Proxy View is of course perfectly coherent, but in contrast to the LLM Theory, it is currently too vague about the theories of HLC that it champions to be of much direct use within empirical science. Empirical science is based on inference to the best explanation and proceeds through competition between reasonably well-understood theories; saying just that one’s favourite theory learns roughly as well as an LLM is insufficient for this purpose. One would like to see enough detail at least for a proper comparison of the relevant theory to generative linguistics’ theories in light of the discoveries and considerations in the literature of the past 70 years or so. More than anything else, the present paper is a plea for supporters of the Proxy View to provide this kind of detail and evaluation for the theory they have in mind.
Still, already at this stage we can outline an empirical evaluation of the Proxy View on various kinds of test cases, and here we focused on two: alignment with the stimulus and cross-linguistic variation. The results were not encouraging for the Proxy View. The LLMs failed to even approximate key pieces of knowledge that humans have, even when the models were trained on much larger corpora than children are exposed to. The LLMs also had an easier time approximating various languages of typologically-unattested kinds than actual human languages, at least with respect to a success criterion used in Kallini et al. (2024). We conclude that, at least as far as the models and phenomena discussed here are concerned, current LLMs do nothing to suggest changes to linguistic theory.
Приложение A Appendix: experiment details
@@ A.1 General Notes
In all APS experiments, probabilities are evaluated on the critical region in boldface (as in the example sentences in Tables A.2, A.2, A.3). For models listed as unidirectional, the probability value takes into consideration the left context of the evaluated token, whereas for bidirectional models, the probability value factors in both the left and the right context. The bidirectional models we consider were pretrained with masked language modelling (fill-mask).
In all generated sentences, the critical region unambiguously disambiguates the sentence as an instance of the phenomenon under consideration (there is no other grammatical reading). To promote lexical variation in target regions, possible target tokens were chosen at random from a POS-tagged English Wikipedia dump. We filtered out potential target tokens that were not part of the vocabulary of one the models considered in our setup, which are detailed in Table 5.
@@
CHILDES Transformer 8.6M Uni Yedetore et al. 2023
BabyLM 10M (strict-small) 10M Bi Warstadt et al. 2023
Wikipedia Transformer 90M Uni Trained by authors
BabyLM 100M 100M Bi Warstadt et al. 2023
BERT Base Uncased $\approx$3.5B Bi Devlin et al. 2018
GPT-2 $\approx$8B Uni Radford et al. 2019
LLaMA 3.2-3b $\approx$9T Uni Meta AI 2024
Таблица 5: Details of the models introduced in Table 2 and used in the experiments in Section 2.
@@ A.2 ATB and PG
For the ATB and PG experiments, we replicate the experimental setup from Lan et al. 2024b with a wider variety of test sentences to increase robustness. We also unify the templates for both phenomena so that they use the same lexical choices where relevant (e.g., proper names). From each template, represented by a context free grammar, we generated tuples of sentences that correspond to the (+filler,+gap), (+filler,-gap) conditions (see CFGs and example sentences in Tables A.2 and A.2). We then evaluated the success criterion P(+filler,gap)<P(+filler,+gap)P(+filler,-gap)<P(+filler,+gap) on a sample of 10,000 tuples for each phenomenon. Underlined words alternate according to the $\pm filler$ condition; words marked in boldface in Tables A.2 and A.2 denote the position where the $\pm gap$ condition becomes evident and probability is measured. We evaluate the performance of all eight models from Table 5, including bidirectional ones. To render the Lan et al. setup viable for testing bidirectional sentences we added a spillover region at the end of every sentence, following a reviewer’s suggestion.
@@
$S\rightarrow\langle{PREAMBLE}\rangle\ \langle{F}\rangle\ \langle{\pm G}\rangle\ \ \langle{SPILLOVER}\rangle\$
$\langle PREAMBLE\rangle\rightarrow\ \textit{I know}\ \langle{WH}\rangle\ \langle{EMBEDDER}\rangle\$
$\langle{WH}\rangle\ \rightarrow\text{`which boys'}\ |\ \text{`which girls'}$
$\langle{EMBEDDER}\rangle\ \rightarrow\text{`you think'}\ |\ \text{`we believe'}$
$\langle{+F}\rangle\ \rightarrow\ \text{that}\ \langle{NAME1}\rangle\ \langle{VP1}\rangle\ \langle{ADV1}\rangle\$
$\langle{+G}\rangle\ \rightarrow\langle{LINK}\rangle\ \ \langle{NAME2}\rangle\ \langle{VP2}\rangle\ \langle{\textbf{ADV2}}\rangle\$
$\langle{-G}\rangle\ \rightarrow\langle{LINK}\rangle\ \ \langle{NAME2}\rangle\ \langle{VP2}\rangle\ \ \langle{\textbf{OBJ}}\rangle\ \ \langle{ADV2}\rangle\$
$\langle{NAME1}\rangle\ \rightarrow\text{`Bob'}\ |\ \text{`John'}$
$\langle{NAME2}\rangle\ \rightarrow\text{`Mary'}\ |\ \text{`Jennifer'}$
$\langle{LINK}\rangle\ \rightarrow\text{`and that'}$
$\langle{ADV1}\rangle\ \rightarrow\text{`shortly'}\ |\ \text{`eventually'}$
$\langle{ADV2}\rangle\ \rightarrow\text{`soon'}\ |\ \text{`today'}\ |\ \text{`now'}$
$\langle{VP1}\rangle\ \rightarrow\text{`will meet'}\ |\ \text{`will see'}$
$\langle{VP2}\rangle\ \rightarrow\ \text{`will hug'}\ |\ \text{`will slap'}\ |\ \text{`will kiss'}$
$\langle{OBJ}\rangle\ \rightarrow\text{`you'}\ |\ \text{`us'}\ |\ \text{`Kim'}$
$\langle{SPILLOVER}\rangle\ \rightarrow\text{`or some other time' | `or next week'}$
$\cdots$
$\Rightarrow$ I know which boys you think that John will meet shortly and that Mary will hug soon or some other time. (+filler,+gap)
$\Rightarrow$ *I know which boys you think that John will meet shortly and that Mary will hug you soon or some other time. (+filler,-gap)
Таблица 6: Excerpt from the context-free grammar used to generate Across-the-Board paradigms for the experiments in Section 2. Critical regions are indicated in boldface. Spillover regions relevant only for bidirectional models are indicated in italics.
PG Grammar
$S\rightarrow\langle{PREAMBLE}\rangle\ \langle{\pm F}\rangle\ \langle{\pm G}\rangle\ \langle{SPILLOVER}\rangle\$
@@
$\langle{+G}\rangle\ \rightarrow\langle{LINK}\rangle\ \langle{V}\rangle\ \langle{\textbf{ADV}}\rangle\$
$\langle{-G}\rangle\ \rightarrow\langle{LINK}\rangle\ \langle{V}\rangle\ \langle{\textbf{OBJ}}\rangle\ \langle{ADV}\rangle\$
$\langle{GEN}\rangle\ \rightarrow\ \textit{'s}$
$\langle{NP}\rangle\ \rightarrow\langle{NP\_SIMPLE}\rangle\ \ |\ \langle{NP\_COMPLEX}\rangle\$
$\langle{NP\_SIMPLE}\rangle\ \rightarrow\langle{GERUND}\rangle\$
$\langle{NP\_COMPLEX}\rangle\ \rightarrow\langle{N\_EMBEDDED}\rangle\ \ \text{`to'}\ \langle{V\_EMBEDDED}\rangle\$
@@
$\langle{V}\rangle\ \rightarrow\text{`bother'}\ |\ \text{`annoy'}\ |\ \text{`disturb'}$
$\langle{OBJ}\rangle\ \rightarrow\text{`you'}\ |\ \text{`us'}\ |\ \text{`Kim'}$
$\langle{GERUND}\rangle\ \rightarrow\text{`talking to'}\ |\ \text{`dancing with'}\ |\ \text{`playing with'}$
$\langle{N\_EMBEDDED}\rangle\ \rightarrow\text{`decision'}\ |\ \text{`intent'}\ |\ \text{`effort'}\ |\ \text{`attempt'}\ |\ \text{`failure'}$
$\langle{V\_EMBEDDED}\rangle\ \rightarrow\text{`talk to'}\ |\ \text{`call'}\ |\ \text{`meet'}\ |\ \text{`dance with'}\ |\ \text{`play with'}$
$\langle{ADV}\rangle\ \rightarrow\text{`soon'}\ |\ \text{`eventually'}$
$\langle{SPILLOVER}\rangle\ \rightarrow\text{'or some other time' | 'or next week'}$
@@
$\Rightarrow$ I know who John’s talking to is going to annoy soon or some other time. (+filler,+gap)
$\Rightarrow$ * I know who John’s talking to is going to annoy you soon or some other time. (+filler,-gap)
Таблица 7: Excerpt from the context-free grammar used to generate Parasitic Gap paradigms for the experiments in Section 2. Critical regions are indicated in boldface. Spillover regions relevant only for bidirectional models are indicated in italics.
@@ A.3 TTE
To create test sets, we generate 1,178,496 sentence pairs, out of which we randomly sample 10,000 pairs per experiment. Paradigms are varied lexically in several ways: we first create paradigms that differ structurally from each other, to control for the possibility that our results are an artifact of the specific sentence structure introduced in the example sentences in Table A.3. We also introduce lexical variation to all non-target elements, such as the main subject, the embedding verb, the wh-element and the verb inside the wh-clause. Most importantly, lexical choices for target verbs and nouns are varied using 30 target nouns and 30 target verbs. We perform five inference experiments on five such samples and consider average accuracy over all experiments. Pairs are evaluated with the success criterion introduced in Section 2.3.
TTE Grammar
$S\rightarrow\langle{PREAMBLE}\rangle\ \langle{F}\rangle\ \langle{EMBEDDING\_VERB}\rangle\ \langle{COMP}\rangle\ \langle{TARGET\_SUBJ}\rangle\ \langle{TARGET\_VERB}\rangle\$
$\langle{PREAMBLE}\rangle\ \rightarrow\text{'He knows'}\ |\ \text{'The boy asked'}\ |\ \text{'The girl knew'}\ldots$
$\langle{F}\rangle\ \rightarrow\text{who}\$
$\langle{EMBEDDING\_VERB}\rangle\ \rightarrow\text{`think'}\ |\ \text{`believe'}\ |\ \text{`say'}$
$\langle{COMP}\rangle\ \rightarrow\text{that}\ |\ \varepsilon$
$\langle{TARGET\_SUBJ}\rangle\ \rightarrow\text{`people'}\ |\ \text{`children'}\ |\ \text{`parents'}\ |\ \text{`scientists'}\ |\ \text{`kids'}\ldots$
$\langle{TARGET\_VERB}\rangle\ \rightarrow\text{'are'}\ |\ \text{'have'}\ |\ \text{'ate'}\ |\ \text{'got'}\ |\ \text{'saw'}\ |\ \text{'met'}\ldots$
$\cdots$
$\Rightarrow$ * He knows who you think that have children (+that,+trace)
$\Rightarrow$ He knows who you think that children have (+that,-trace)
$\Rightarrow$ He knows who you think have children (-that,+trace)
$\Rightarrow$ He knows who you think children have (-that,-trace)
Таблица 8: Excerpt from the context-free grammar used to generate That-Trace-Effect paradigms for the experiments in Section 2.
Приложение B Appendix: datasets and model building